# Toward Universal Decoding of Binary Linear Block Codes via Enhanced Polar Transformations

### v1

@@ III-A2 Multi-kernel technique to meet the code length
In ${\mathsf{PD}}$, the multi-kernel technique was adopted to decode BLBC whose code lengths are not powers of 2. For example, when decoding the $(24,12)$ extended Golay code, [14] considers $\mathbf{G}_{24}=\mathbf{F}_{3}\otimes\mathbf{F}_{2}^{\otimes 3}$, where
$F_{3}=\left[\begin{array}[]{cccccccc}1&1&0\\ 1&0&1\\ 1&1&1\end{array}\right].$ (16)
@@ III-A3 Search algorithm
Motivated by the above drawbacks, we proposed ${\mathsf{PD}^{+}}$, which addresses each of these drawbacks.
@@ III-B Main Result and Proposed ${\mathsf{PD}^{+}}$
Now, we present the main result: An enhanced version of the ${\mathsf{PD}}$ decoding algorithm from [14]. In ${\mathsf{PD}^{+}}$, to decode an $(n,k)$ code with a $k\times n$ generator matrix $\mathbf{G}$, we begin with the $N\times N$ generator matrix $\mathbf{G}_{N}$ of the traditional polar code of size $N\geq n$, using Arikan’s kernel as defined in (4). We then prune the connections in $\mathbf{G}_{N}$ according to the $N\times\log_{2}(N)$ pruning matrix $\mathbf{R}$, resulting in ${\tilde{\mathbf{G}}_{N}}=f(\mathbf{R},\mathbf{G}_{N})$ the generator matrix of a pruned polar code $\tilde{\mathcal{C}}_{p}$. The details of the pruning pattern $\mathbf{R}$ and the pruning function $f$ will be discussed in Section III-C. We now state the following theorem:
@@
For an $(n,k)$ BLBC $\mathcal{C}$, an integer $N>n$, an $N\times N$ permutation matrix $\mathbf{P}$, and a $N\times n$ shortening matrix $\mathbf{S}$ containing $n$ rows of the identity matrix $\mathbf{I}_{N}$, there exists a polar-like code [^5] Here, a polar-like code refers to a shortened version of a pruned polar code. $\tilde{\mathcal{C}}_{p}$ with dynamic frozen bits, such that the one-to-one correspondence between the codeword $\mathbf{c}$ in $\mathcal{C}$ and the codeword $\mathbf{c}_{p}$ in $\tilde{\mathcal{C}}_{p}$ is given by $\mathbf{c}=\mathbf{c}_{p}\mathbf{P}\mathbf{S}$.
@@
$\mathcal{C}\triangleq$$\left\{\mathbf{c}=\mathbf{mG}|\mathbf{m}\in\mathbb{F}_{2}^{k}\right\}=\left\{\mathbf{c}=\mathbf{m}\mathbf{I}_{k}\mathbf{G}\mathbf{I}_{N}|\mathbf{m}\in\mathbb{F}_{2}^{k}\right\}$
$=$ $\left\{\mathbf{c}=\mathbf{m}(\mathbf{E}^{-1}\mathbf{E})\mathbf{G}(\mathbf{S}^{{\dagger}}\mathbf{P}^{-1}{\tilde{\mathbf{G}}_{N}}^{-1}{\tilde{\mathbf{G}}_{N}}\mathbf{P}\mathbf{S})|\mathbf{m}\in\mathbb{F}_{2}^{k}\right\}$
$=$ $\left\{\mathbf{c}=(\mathbf{m}\mathbf{E}^{-1})(\mathbf{E}\mathbf{G}\mathbf{S}^{{\dagger}}\mathbf{P}^{-1}{\tilde{\mathbf{G}}_{N}}^{-1}){\tilde{\mathbf{G}}_{N}}\mathbf{P}\mathbf{S}|\mathbf{m}\in\mathbb{F}_{2}^{k}\right\}$
$=$ $\left\{\mathbf{c}=\mathbf{m}_{p}\mathbf{M}_{\rm DF}{\tilde{\mathbf{G}}_{N}}\mathbf{P}\mathbf{S}|\mathbf{m}_{p}\in\mathbb{F}_{2}^{k}\right\}$
$=$ $\left\{\mathbf{c}=\mathbf{c}_{p}\mathbf{P}\mathbf{S}|\mathbf{c}_{p}\in\tilde{\mathcal{C}}_{p}\right\},$ (17)
@@
and interpret it as the received signal corresponding to the output of the channel with input from $\tilde{\mathcal{C}}_{p}$. Note that although $\tilde{\mathcal{C}}_{p}$ is pruned, its generator matrix ${\tilde{\mathbf{G}}_{N}}$ retains the structure of Arikan’s polar code $\mathbf{G}_{N}$, enabling decoding by polar code algorithms with minor modifications to be discussed in Section III-C. A block diagram summarizing the decoding process is provided in Fig 3.
@@
We would like to emphasize that $\mathbf{P}^{-1}$, $\mathbf{S}^{\dagger}$, and $\mathbf{R}$ are all precomputed offline in advance. Consequently, the proposed ${\mathsf{PD}^{+}}$ merely applies these operations. Moreover, $\mathbf{P}^{-1}$ is a permutation matrix, $\mathbf{S}^{\dagger}$ simply adds 0 back, and $\mathbf{R}$ just disables some edges in $\mathbf{G}_{N}$, resulting in minimal additional complexity. ∎
@@ III-C Pruned Kernel
To represent the pruning pattern, we use an $N/2\times\log_{2}(N)$ binary matrix $\mathbf{R}$, where a value of 1 indicates pruning the edge, and 0 indicates no pruning. Additionally, we denote the pruning operation by $f$ and the generator matrix of the corresponding pruned polar code $\tilde{\mathcal{C}}_{p}$ by ${\tilde{\mathbf{G}}_{N}}=f(\mathbf{R},\mathbf{G}_{N})$. An example of the pruned polar code is provided in Example 4.
@@
Consider the polar code with $N=8$ constructed with Arikan’s kernel as shown in Fig. 4. In this example, we prune 2 edges in the first stage, 1 edge in the second stage, and 1 edge in the third stage. The corresponding pruning matrix is given by
$\mathbf{R}=\left[\begin{array}[]{ccc}\color[rgb]{1,0,0}0&1&1\\[-4.55244pt] 1&1&\color[rgb]{0,1,0}0\\[-4.55244pt] 1&\color[rgb]{0,0,1}0&1\\[-4.55244pt] \mathbf{0}&1&1\end{array}\right],$ (19)
@@
It is worth emphasizing that this pruned kernel retains the butterfly-like structure, ensuring compatibility with both SC and SCL decoders. When a connection is pruned, the data passes through without undergoing polarization. Initially, we have two operations: ${\mathcal{W}}^{-}={\mathcal{W}}_{a}\boxplus{\mathcal{W}}_{b}$ and ${\mathcal{W}}^{+}={\mathcal{W}}_{a}\boxtimes{\mathcal{W}}_{b}$. For the pruned kernel, however, we introduce an additional operation, ${\mathcal{W}}_{a}\square{\mathcal{W}}_{b}$, defined by the following relationships:
$({\mathcal{W}}_{a}\square{\mathcal{W}}_{b})(y_{1},y_{2}|u_{1})\triangleq{\mathcal{W}}_{a}(y_{1}|u_{1}),$ (20)
@@
$({\mathcal{W}}_{a}\square{\mathcal{W}}_{b})(y_{1},y_{2},u_{1}|u_{2})\triangleq{\mathcal{W}}_{b}(y_{2}|u_{2}).$ (21)
@@
Polar codes with pruned kernels were previously investigated in [32], where a kernel is pruned if the two incoming bits are either both linear combinations of message bits or both linear combinations of frozen bits. Consequently, the pruning pattern in the latter stages influences the patterns in earlier stages. Using this approach, the authors of [32] were able to construct polar codes with fixed frozen bits that achieve improved error exponents.
In contrast, the present work aims to transform a BLBC into a polar-like code with dynamic frozen bits. Imposing such restrictions on pruning patterns would significantly limit the flexibility of our proposed ${\mathsf{PD}^{+}}$. Therefore, we allow all $N/2\cdot\log_{2}(N)$ pruning patterns to be used. ∎
@@ III-D Shortening Operation
In this remark, we emphasize that the proposed pruning and shortening operations together enable the generation of a diverse set of kernels. To illustrate this, we consider the example in Fig. 5. In this example, we begin with a size-4 polar code constructed from $\mathbf{F}_{2}\otimes\mathbf{F}_{2}$, demonstrating that all the 8 distinct lower-triangluar kernels of size 3 can be derived through shortening and puncturing. This phenomenon is general, and the diversity of kernels can be significantly expanded by leveraging pruning and shortening, even though we always start with Arikan’s kernels. ∎
Figure 5: Example of using pruning and shortening to get kernels of size 3 from the Arikan’s kernel.
@@ IV AI-Inspired Searching Algorithm
According to the discussion in Section III, the proposed method involves three parameters that must be optimized: 1) the permutation operation governed by the permutation matrix $\mathbf{P}$, 2) the pruning operation governed by the pruning function identified as $\mathbf{R}$, and 3) the shortening operation governed by the shortening indices $\mathcal{S}$. Before introducing the AI-inspired search algorithm, we first demonstrate that the searches for the permutation and shortening matrices can be combined to reduce complexity.
@@
For $k\in{1,2}$, the shortening operation simply excludes some indices $\mathcal{S}_{k}\subseteq\{1,\ldots,N\}$ of $\mathbf{c}$. This can be expressed as first forming $\mathbf{c}\mathbf{S}_{k}$, where $\mathbf{S}_{k}$ is constructed by setting the columns corresponding to $\mathcal{S}_{k}$ in the identity matrix $\mathbf{I}_{N}$ to $\mathbf{0}$, and then transmitting only the indices in $\mathcal{S}_{k}^{c}$. Since $\mathcal{S}_{k}$ is known at both the sender and receiver, the receiver can pad zeros back into the excluded positions and perform decoding as if no shortening had occurred. Given that $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|$, it follows that there exists a permutation matrix $\mathbf{T}$ such that $\mathbf{S}_{1}\mathbf{T}=\mathbf{S}_{2}$. ∎
@@
In Section III, we have identified the pruning pattern as a $\frac{N}{2}\times\log_{2}{N}$ binary matrix $\mathbf{R}$; each bit acts as a flag to denote a connection status. The permutation matrix can be identified as a length-$N$ integer vector $\mathbf{p}$, where the $i$-th number $p_{i}$ indicates that we will reorder the $i$-th bit to the $p_{i}$-th bit. An example is given in what follows:
@@
6:     Randomly generate $(\mathbf{R}_{\text{next}},\mathbf{p}_{\text{next}})$ by either flipping a single bit of $\mathbf{R}_{\text{current}}$ or swapping two elements of $\mathbf{p}_{\text{current}}$
7:     Calculate $p_{e,\text{next}}$ using (13) with input $(\mathbf{R}_{\text{next}},\mathbf{p}_{\text{next}})$
8:     if  $p_{e,\text{next}}<p_{e,\text{current}}$  then
9:         $(\mathbf{R}_{\text{current}},\mathbf{p}_{\text{current}})\leftarrow(\mathbf{R}_{\text{next}},\mathbf{p}_{\text{next}})$ always
10:     else
11:         $(\mathbf{R}_{\text{current}},\mathbf{p}_{\text{current}})\leftarrow(\mathbf{R}_{\text{next}},\mathbf{p}_{\text{next}})$ with probability $\exp(-(p_{e,\text{next}}-p_{e,\text{current}})/T)$
12:     end if
13:end while
14:return $(\mathbf{R}_{\text{current}},\mathbf{p}_{\text{current}})$
@@ V Simulation Results
In this section, we present simulation results to validate the effectiveness of the proposed ${\mathsf{PD}^{+}}$. For comparison, we also include the performance of several benchmark decoding algorithms: OSD, GRAND, [^7] For GRAND, we implement the segmented ORB-GRAND in [33] by adapting the computer programs in https://github.com/mohammad-rowshan/Segmented-GRAND. ${\mathsf{PD}}$, and, where feasible, MLD. In the legend, OSD represents OSD of order $\ell$, while GRANDM denotes GRAND with $M$ candidate noise sequences. For ${\mathsf{PD}}$ and ${\mathsf{PD}^{+}}$, the subscripts indicate their respective list sizes in SCL decoding. Beyond performance evaluation, we also compare the computational complexity of the considered schemes. Following the approach in [give a reference here], complexity is measured by the number of candidate sequences the decoder processes.
In Section V-A, the challenging case shown in (1) is addressed. In Section V-B, extended BCH codes are considered. In Section V-C is revisited, which is followed by simulations for binary quadratic residue codes in Section V-D. [^8] All the $(\mathbf{R},\mathbf{P})$ pairs used in our simulations are available at the following link: https://github.com/thisistt/Enhanced-polar-decoding.

### v2

@@ III-A2 Multi-kernel technique to meet the code length
In ${\mathsf{PD}}$, the multi-kernel technique was adopted to decode BLBC whose code lengths are not powers of 2. For example, when decoding the $(24,12)$ eGolay code, [14] considers $\mathbf{G}_{24}=\mathbf{F}_{3}\otimes\mathbf{F}_{2}^{\otimes 3}$, where
$\mathbf{F}_{3}=\left[\begin{array}[]{cccccccc}1&1&0\\ 1&0&1\\ 1&1&1\end{array}\right].$ (16)
@@ III-A3 Search algorithm
Motivated by the above drawbacks, we propose ${\mathsf{PD}^{+}}$, which addresses each of these drawbacks.
@@ III-B Main Result and Proposed ${\mathsf{PD}^{+}}$
Now, we present the main result: An enhanced version of the ${\mathsf{PD}}$ decoding algorithm from [14]. In ${\mathsf{PD}^{+}}$, to decode an $(n,k)$ code with a $k\times n$ generator matrix $\mathbf{G}$, we begin with the $N\times N$ generator matrix $\mathbf{G}_{N}$ of the traditional polar code of size $N\geq n$, using Arıkan’s kernel as defined in (4). We then prune the connections in $\mathbf{G}_{N}$ according to the $N\times\log_{2}(N)$ pruning matrix $\mathbf{R}$, resulting in ${\tilde{\mathbf{G}}_{N}}=f(\mathbf{R},\mathbf{G}_{N})$ the generator matrix of a pruned polar code $\tilde{\mathcal{C}}_{p}$. The details of the pruning pattern $\mathbf{R}$ and the pruning function $f$ will be discussed in Section III-C. We now state the following theorem:
@@
For an $(n,k)$-BLBC $\mathcal{C}$, an integer $N>n$, an $N\times N$ permutation matrix $\mathbf{P}$, and an $N\times n$ shortening matrix $\mathbf{S}$ containing $n$ columns of the identity matrix $\mathbf{I}_{N}$, there exists a polar-like code [^4] Here, a polar-like code refers to a shortened version of a pruned polar code. $\tilde{\mathcal{C}}_{p}$ with dynamic frozen bits, such that the one-to-one correspondence between the codeword $\mathbf{c}$ in $\mathcal{C}$ and the codeword $\mathbf{c}_{p}$ in $\tilde{\mathcal{C}}_{p}$ is given by $\mathbf{c}=\mathbf{c}_{p}\mathbf{P}\mathbf{S}$.
@@
$\mathcal{C}\triangleq$$\left\{\mathbf{c}=\mathbf{mG}\mid\mathbf{m}\in\mathbb{F}_{2}^{k}\right\}=\left\{\mathbf{c}=\mathbf{m}\mathbf{I}_{k}\mathbf{G}\mathbf{I}_{N}\mid\mathbf{m}\in\mathbb{F}_{2}^{k}\right\}$
$=$ $\left\{\mathbf{c}=\mathbf{m}(\mathbf{E}^{-1}\mathbf{E})\mathbf{G}(\mathbf{S}^{{\dagger}}\mathbf{P}^{-1}{\tilde{\mathbf{G}}_{N}}^{-1}{\tilde{\mathbf{G}}_{N}}\mathbf{P}\mathbf{S})\mid\mathbf{m}\in\mathbb{F}_{2}^{k}\right\}$
$=$ $\left\{\mathbf{c}=(\mathbf{m}\mathbf{E}^{-1})(\mathbf{E}\mathbf{G}\mathbf{S}^{{\dagger}}\mathbf{P}^{-1}{\tilde{\mathbf{G}}_{N}}^{-1}){\tilde{\mathbf{G}}_{N}}\mathbf{P}\mathbf{S}\mid\mathbf{m}\in\mathbb{F}_{2}^{k}\right\}$
$=$ $\left\{\mathbf{c}=\mathbf{m}_{p}\mathbf{M}_{\rm DF}{\tilde{\mathbf{G}}_{N}}\mathbf{P}\mathbf{S}\mid\mathbf{m}_{p}\in\mathbb{F}_{2}^{k}\right\}$
$=$ $\left\{\mathbf{c}=\mathbf{c}_{p}\mathbf{P}\mathbf{S}\mid\mathbf{c}_{p}\in\tilde{\mathcal{C}}_{p}\right\},$ (17)
@@
and interpret it as the received signal corresponding to the output of the channel with input from $\tilde{\mathcal{C}}_{p}$. Note that although $\tilde{\mathcal{C}}_{p}$ is pruned, its generator matrix ${\tilde{\mathbf{G}}_{N}}$ retains the structure of Arıkan’s polar code $\mathbf{G}_{N}$, enabling decoding by polar code algorithms with minor modifications to be discussed in Section III-C. A block diagram summarizing the decoding process is provided in Fig 3.
@@
We would like to emphasize that $\mathbf{P}^{-1}$, $\mathbf{S}^{\dagger}$, and $\mathbf{R}$ are all precomputed offline in advance. Consequently, the proposed ${\mathsf{PD}^{+}}$ merely applies these operations. Moreover, $\mathbf{P}^{-1}$ is a permutation matrix, $\mathbf{S}^{\dagger}$ simply adds 0 back, [^5] Since $\mathbf{S}$ is composed of columns of $\mathbf{I}_{N}$, $\mathbf{S}^{\dagger}$ is simply $\mathbf{S}^{T}$. In (17), we write $\mathbf{S}^{\dagger}$ instead of $\mathbf{S}^{T}$ to emphasize its role in inverting the operation of $\mathbf{S}$. and $\mathbf{R}$ just disables some edges in $\mathbf{G}_{N}$, resulting in minimal additional complexity. ∎
@@ III-C Pruned Kernel
To represent the pruning pattern, we use an $N/2\times\log_{2}(N)$ binary matrix $\mathbf{R}$, where a value of 0 indicates pruning the edge, and 1 indicates no pruning. Additionally, we denote the pruning operation by $f$ and the generator matrix of the corresponding pruned polar code $\tilde{\mathcal{C}}_{p}$ by ${\tilde{\mathbf{G}}_{N}}=f(\mathbf{R},\mathbf{G}_{N})$. An example of the pruned polar code is provided in Example 4.
@@
Consider the polar code with $N=8$ constructed with Arıkan’s kernel as shown in Fig. 4. In this example, we prune 2 edges in the first stage, 1 edge in the second stage, and 1 edge in the third stage. The corresponding pruning matrix is given by
$\mathbf{R}=\left[\begin{array}[]{ccc}\color[rgb]{1,0,0}0&1&1\\[-4.55244pt] 1&1&\color[rgb]{0,1,0}0\\[-4.55244pt] 1&\color[rgb]{0,0,0}0&1\\[-4.55244pt] \mathbf{0}&1&1\end{array}\right],$ (19)
@@
[…]
$({\mathcal{W}}_{a}\mathbin{\square}{\mathcal{W}}_{b})(y_{1},y_{2}|u_{1})\triangleq{\mathcal{W}}_{a}(y_{1}|u_{1}),$ (20)
@@
$({\mathcal{W}}_{a}\mathbin{\square}{\mathcal{W}}_{b})(y_{1},y_{2},u_{1}|u_{2})\triangleq{\mathcal{W}}_{b}(y_{2}|u_{2}).$ (21)
@@
Prior work has explored polar codes with pruned kernels. The intuition has been that pruning Arıkan kernels inherently degrades polar code performance by reducing the mixing crucial for channel polarization. Consequently, pruning has primarily served as a technique to decrease decoding complexity. Take [32] for example, where a kernel is pruned if the two incoming bits are either both linear combinations of message bits or both linear combinations of frozen bits. Consequently, the pruning pattern in the latter stages influences the patterns in earlier stages. Using this idea, the authors of [32] were able to come up with a pruning strategy that constructs polar codes that achieve the capacity for binary erasure channels, while enjoying $\log\log(N)$ per-bit complexity.
In contrast, unlike the construction of standard polar codes, where the selection of the frozen set is a design choice, our problem is dictated by the inherent structure of $(n,k)$-BLBC. Moreover, since $n$ may not always be a power of 2, we are naturally led to consider non-Arıkan kernels with sizes that are also not powers of 2. Pruning, in conjunction with shortening, offers a means to generate non-standard kernels with the desired dimensions as illustrated in Figs. 4 and 5. Therefore, the intuition that pruning might negatively impact SC decoding performance may not directly apply to our fundamentally different objective of decoding an $(n,k)$-BLBC.
@@ III-D Shortening Operation
In this remark, we emphasize that the proposed pruning and shortening operations together enable the generation of a diverse set of kernels. To illustrate this, we consider the example in Fig. 5. In this example, we begin with a size-4 polar code constructed from $\mathbf{F}_{2}\otimes\mathbf{F}_{2}$, demonstrating that all the 8 distinct lower-triangluar kernels of size 3 can be derived through shortening and puncturing. This phenomenon is general, and the diversity of kernels can be significantly expanded by leveraging pruning and shortening, even though we always start with Arıkan’s kernels. ∎
Figure 5: Example of using pruning and shortening to get kernels of size 3 from the Arıkan’s kernel.
@@ IV AI-Inspired Searching Algorithm
According to the discussion in Section III, the proposed method involves three parameters that must be optimized: 1) the permutation operation governed by the permutation matrix $\mathbf{P}$, 2) the pruning operation governed by the pruning function identified as $\mathbf{R}$, and 3) the shortening operation governed by the shortening indices $\mathcal{S}$. The goal is to find $\mathbf{P}$, $\mathbf{R}$, and $\mathcal{S}$ such that the sum of the Bhattacharyya parameters in (13) is minimized. Although this sum serves as an upper bound on SC decoding performance, it is often regarded as a good indicator of the performance of more sophisticated decoders (see, for example, [33, Conjecture 1]).
Before introducing the AI-inspired search algorithm, we first demonstrate that the searches for the permutation and shortening matrices can be combined to reduce complexity.
@@
For $k\in\{1,2\}$, the shortening operation simply excludes some indices $\mathcal{S}_{k}\subseteq\{1,\ldots,N\}$ of $\mathbf{c}$. This can be expressed as first forming $\mathbf{c}\mathbf{S}_{k}$, where $\mathbf{S}_{k}$ is constructed by setting the columns corresponding to $\mathcal{S}_{k}$ in the identity matrix $\mathbf{I}_{N}$ to $\mathbf{0}$, and then transmitting only the indices in $\mathcal{S}_{k}^{c}$. Since $\mathcal{S}_{k}$ is known at both the sender and receiver, the receiver can pad zeros back into the excluded positions and perform decoding as if no shortening had occurred. Given that $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|$, it follows that there exists a permutation matrix $\mathbf{T}$ such that $\mathbf{S}_{1}\mathbf{T}=\mathbf{S}_{2}$. ∎
@@
In Section III, we have identified the pruning pattern as a $\frac{N}{2}\times\log_{2}{N}$ binary matrix $\mathbf{R}$; each bit acts as a flag to denote a connection status. The permutation matrix can be identified as an one-line notation with length-$N$ integer vector $\mathbf{p}$, where the $i$-th number $p_{i}$ indicates that we will reorder the $i$-th bit to the $p_{i}$-th bit. An example is given in what follows:
@@
6:     Randomly generate $(\mathbf{R}_{\text{next}},\mathbf{p}_{\text{next}})$ by either flipping a
7:     single bit of $\mathbf{R}_{\text{current}}$ or swapping two elements of
8:     $\mathbf{p}_{\text{current}}$
9:     Calculate $p_{e,\text{next}}$ using (13) with input $(\mathbf{R}_{\text{next}},\mathbf{p}_{\text{next}})$
10:     if  $p_{e,\text{next}}<p_{e,\text{current}}$  then
11:         $(\mathbf{R}_{\text{current}},\mathbf{p}_{\text{current}})\leftarrow(\mathbf{R}_{\text{next}},\mathbf{p}_{\text{next}})$ always
12:     else
13:         $(\mathbf{R}_{\text{current}},\mathbf{p}_{\text{current}})\leftarrow(\mathbf{R}_{\text{next}},\mathbf{p}_{\text{next}})$ with probability
14:          $\exp(-(p_{e,\text{next}}-p_{e,\text{current}})/T)$
15:     end if
16:end while
17:return $(\mathbf{R}_{\text{current}},\mathbf{p}_{\text{current}})$
@@ V Simulation Results
In this section, we present simulation results to validate the effectiveness of the proposed ${\mathsf{PD}^{+}}$. For comparison, we also include the performance of several benchmark decoding algorithms: OSD, GRAND, [^7] For GRAND, we implement the segmented ORB-GRAND in [34] by adapting the computer programs in https://github.com/mohammad-rowshan/Segmented-GRAND. ${\mathsf{PD}}$, and, where feasible, MLD. In the legend, OSD represents OSD of order $\ell$, while GRANDM denotes GRAND with $M$ candidate noise sequences. For ${\mathsf{PD}}$ and ${\mathsf{PD}^{+}}$, the subscripts indicate their respective list sizes in SCL decoding. Beyond performance evaluation, we also compare the computational complexity of the considered schemes. Following the approach in [12, 34], complexity is measured by the number of candidate sequences the decoder processes.
In Section V-A, the challenging case shown in (1) is addressed. In Section V-B, eBCH codes are considered. In Section V-C is revisited, which is followed by simulations for binary quadratic residue codes in Section V-D. Last but not least, in Section V-E, we present simulation results for a randomly generated code to address the potential concern that the proposed ${\mathsf{PD}^{+}}$ is only effective for codes like eBCH, eGolay, and binary QR codes, which possess very strong algebraic structures. [^8] All the $(\mathbf{R},\mathbf{P})$ pairs used in our simulations are available at the following link: https://github.com/thisistt/Enhanced-polar-decoding.